Let ?(?)y(t) be the solution to ?′=?+?y′=t+y satisfying ?(5)=6.satisfying y(5)=6.

Use Euler's Method with time step ℎ=0.1h=0.1 to approximate ?(5.5).approximate y(5.5).

(Use decimal notation. Give your answers to four decimal places.)n= 0, to = 5, yo = n = 1, 11 = 5.1, yı = n = 2, 12 = 5.2, y2 = n = 3,13 = 5.3, y3 = n = 4, 14 = 5.4, y4 = n = 5, t5 = 5.5, y5
Answer #1

1-0. to = 5,426 (64)=(+9 ti ७) = 90 + h Ce.go) 6+ 7.7 36 = 5.२ फर y, tht (ti, g) म. + ! - (२.२ 10 8.32 t3 = 5.3 93- 8.32+ - X